OpenStudy (anonymous):

15(2 - x) = 13(3 - x)

OpenStudy (anonymous):

30-15x=39-13x -9=2x x=-9/2

OpenStudy (anonymous):

30-15x = 39 -13x 30-39 = -13x + 15x -9 = 2x -9/2 = x